OpenStudy (ashleynguyenx3):

Write the Maclaurin Series (including the general term) for f(x)=ln(e+x).

mathslover (mathslover):

Hmn do you know what is Maclaurin series?

OpenStudy (ashleynguyenx3):

Yes, I do

mathslover (mathslover):

Ok! If I explain you how to get the MacLaurin series for f(x) = ln(1+x) , will you be able to get the Maclaurin series for f(x) = ln(e+x) ?

OpenStudy (ashleynguyenx3):

Yes, since e is just a constant

mathslover (mathslover):

Ok so our new question is : Write the Maclaurin series for f(x) = ln (1+x) .

mathslover (mathslover):

First of all, the center of our Taylor series will be a = 0 . So , let us construct a table that will give us derivatives and enable us to calculate \(f^{(n)}(a)\)

OpenStudy (ashleynguyenx3):

For ln(1+x)

mathslover (mathslover):

Please wait, it will take time for me to create the table.

OpenStudy (ashleynguyenx3):

Okay :D

mathslover (mathslover):

OS does not support table latex here :(

OpenStudy (ashleynguyenx3):

Aww, have any other way?

mathslover (mathslover):

Yes , please wait for 1minute more.

OpenStudy (ashleynguyenx3):

Use print screen?

mathslover (mathslover):

Sorry

OpenStudy (ashleynguyenx3):

It's fine

mathslover (mathslover):

\(\begin{array}{|c|c|c|} \hline \textbf{n} & f^{(n)}(x)& f^{(n)}(a)(0)\\ \hline n & f^{(n)}(x) & f^{(n)}(a)(0) \\ \hline 0 & ln(1+x) & ln(1)= 0 \\ \hline 1 & (1+x)^{-1} & 1 \\ \hline 2 & -(1+x)^{-2} & -1 \\ \hline 3 & +2(1+x)^{-3} & +2 \\ \hline 4 & -2.3(1+x)^{-4} & -2.3 \\ \hline . & . & . \\ \hline . & . & . \\ \hline . & . & . \\ \hline n \ne 0 & (-n)^{n+1}(n-1)! \cfrac{1}{(1+x)^n} & (-1)^{n+1}(n-1)! \\ \hline \end{array}\)

mathslover (mathslover):

@ashleynguyenx3 sorry for being so late. but finally I made it :)

OpenStudy (ashleynguyenx3):

YAY! :D

hero (hero):

Excellent job @mathslover

OpenStudy (ashleynguyenx3):

So it'll basically be the same right?

mathslover (mathslover):

So now the general form is : \(f^{(n)}(2) = (-1)^{n+1}(n-1)! \) if \(n \ne 0\) , and and \(f^{(0)}(0)=0\) . Since the form hanged, we will have to pull the n = 0 term out of our sum .

mathslover (mathslover):

Thank you, hero.

mathslover (mathslover):

\(\large{c_n = \cfrac{f^{(n)}(0)}{n!} = \cfrac{(-1)^{n+1}(n-1)!}{n!} = \cfrac{(-1)^{n+1}}{n} , \space n \ne 0 ; \space c_0 = 0}\)

mathslover (mathslover):

So the taylor series is given by : \(\large{\ln (1+x) = 0 + \sum_{n=1}^{{\infty}} \cfrac{f^{(n)}(0)}{n!}(x-0)^n = \sum_{n=1}^{\infty}\cfrac{(-1)^{n+1}}{n} x^n , |x|< R }\)

mathslover (mathslover):

Can you do it now, @ashleynguyenx3 ?

mathslover (mathslover):

You just have to calculate radius of the convergence, R. You can do it by using the ratio test : \(\large{a_n = \cfrac{(-1)^{n+1}}{n} x^n}\)

OpenStudy (ashleynguyenx3):

Yeah, I know that's the second part to my problem.

mathslover (mathslover):

So you got it?

OpenStudy (ashleynguyenx3):

Um yeah I think

OpenStudy (ashleynguyenx3):

I'm gonna try to solve it right now

mathslover (mathslover):

Good! Let us know what you get, and also if you have any problem , we will try to solve it.

OpenStudy (ashleynguyenx3):

Okays thank you! :D

mathslover (mathslover):

You're welcome ashley. \(\textbf{Welcome!} \space \ddot \smile \)

OpenStudy (anonymous):

Very complete solution mathslover :) do u know how to do this with geometric series? have u learnt integral yet?

OpenStudy (anonymous):

im asking ashley too :)

mathslover (mathslover):

Thank you mukushla , ok you're asking that to me and ashley ?

OpenStudy (ashleynguyenx3):

I thought you use geometric series to just find the convergence or divergence..

mathslover (mathslover):

Well, I have started learning integral already . though, not in soo deep, but thanks to some sites and wikipedia that provided me with some extra knowledge. I saw such a type of question in openstudy too, just saw it and when I scrolled up and down the wikipedia page of MacLaurin series, i remembered everything.

OpenStudy (anonymous):

very well, then i refuse to bring that solution here :) because it needs integral calculus :)

OpenStudy (ashleynguyenx3):

@mathslover I don't know if I did this right..

Parth (parthkohli):

There's a better solution with integrals as @mukushla said. But I'd just have done what mathslover has :-D

OpenStudy (ashleynguyenx3):

How would you use integrals?

mathslover (mathslover):

Friends, please post your opinions too, it will be better if the user learns something easier and better soln for this question. @mukushla @ParthKohli

Parth (parthkohli):

I don't know how to do it exactly, but I read on MSE about solving an integral through Maclaurin Series. I'm pretty sure Mukushla is talking just the vice-versa. But as I said, I would choose what Mathslover did.

Parth (parthkohli):

I guess we can kinda integrate \(\ln(e + x)\), find the Maclaurin Series for that, and then differentiate it.

OpenStudy (ashleynguyenx3):

What's the anti derivative of lnx?

Parth (parthkohli):

\[x(\ln(x) - 1)\]I guess

Parth (parthkohli):

\[+C\]

OpenStudy (ashleynguyenx3):

So it'd be (e+x)(ln(e+x)-1)+c

OpenStudy (anonymous):

so its like this friends, geometric series of \(\frac{1}{1+x}\) for \(|x|<1\)\[\frac{1}{1+x}=1-x+x^2-x^3+...\]take integral of both sides\[\int_{0}^{x} \frac{1}{1+x} \ \text{d}x=\int_{0}^{x} (1-x+x^2-x^3+...) \ \text{d}x\]\[\ln(1+x)=x-\frac{x^2}{2}+\frac{x^3}{3}-\frac{x^4}{4}+...+\frac{(-1)^{n+1}}{n}x^n+...\]

mathslover (mathslover):

This looks something that I have studied somewhere

mathslover (mathslover):

Great work Mukushla too

OpenStudy (anonymous):

;) Tnx :)